# Weight-Entanglement Meets Gradient-Based Neural Architecture Search

### v1

@@ Abstract
Weight sharing is a fundamental concept in neural architecture search (NAS), enabling gradient-based methods to explore cell-based architecture spaces significantly faster than traditional blackbox approaches. In parallel, weight entanglement has emerged as a technique for intricate parameter sharing among architectures within macro-level search spaces. Since weight-entanglement poses compatibility challenges for gradient-based NAS methods, these two paradigms have largely developed independently in parallel sub-communities. This paper aims to bridge the gap between these sub-communities by proposing a novel scheme to adapt gradient-based methods for weight-entangled spaces. This enables us to conduct an in-depth comparative assessment and analysis of the performance of gradient-based NAS in weight-entangled search spaces. Our findings reveal that this integration of weight-entanglement and gradient-based NAS brings forth the various benefits of gradient-based methods (enhanced performance, improved supernet training properties and superior any-time performance), while preserving the memory efficiency of weight-entangled spaces. The code for our work is openly accessible here.
@@ 1 Introduction
The concept of weight-sharing in Neural Architecture Search (NAS) arose from the need to improve the efficiency of conventional blackbox NAS algorithms, which demand significant computational resources to evaluate individual architectures. Here, weight-sharing (WS) refers to the paradigm by which we represent the search space with a single large supernet, also known as the one-shot model, that subsumes all the candidate architectures in that space. Every edge of this supernet holds all the possible operations that can be assigned to that edge. Importantly, architectures that share a particular operation also share its corresponding operation weights, allowing simultaneous training of an exponential number of subnetworks, unlike the sequential approach of blackbox NAS.
@@
Gradient-based NAS algorithms (or optimizers), such as DARTS (Liu et al., 2019), GDAS (Dong and Yang, 2019) and DrNAS (Chen et al., 2021b), assign an architectural parameter to every choice of operation on a given edge of the supernet. The output feature maps of these edges are thus an aggregation of the outputs of the individual operations on that edge, weighted by their architectural parameters. These architectural parameters are learned using gradient updates by differentiating through the validation loss. Supernet weights and architecture parameters are therefore trained simultaneously in a bi-level fashion. Once this training phase is complete, the final architecture can be identified quickly, e.g., by selecting operations with the highest architectural weights on each edge as depicted in Figure 1(b). However, more sophisticated methods have also been explored (Wang et al., 2021) for this selection.
@@
While gradient-based NAS methods have primarily been studied for cell-based NAS search spaces, a different class of search spaces focuses on macro-level structures (parameterizing kernel size, number of channels, etc.) for which all architectures in the space are subnetworks of the architecture with the largest architectural choices, which is identical to the supernet in this case. These search spaces share weights via the more intricate form of weight-entanglement (WE) between similar operations on the same edge; e.g., the nine weights of a 3 $\times$ 3 convolution are a subset of the 25 weights of a 5 $\times$ 5 convolution. This paradigm reduces the memory requirements of the supernet to the size of the largest architecture in the search space.
@@
In order to efficiently search over such weight-entanglement spaces, two-stage methods have been introduced that first pre-train the supernet, and then perform blackbox search on it to obtain the final architecture. OFA (Cai et al., 2020), SPOS (Guo et al., 2020), AutoFormer (Chen et al., 2021a) and HAT (Wang et al., 2020) are prominent examples of methods that fall into this category. Note that these methods do not employ additional architectural parameters for supernet training or search. They typically train the supernet by randomly sampling subnetworks and training them as depicted in Figure 1(a). The post-hoc blackbox search relies on using the performance of subnetworks sampled from the trained supernet as a proxy for true performance on the unseen test set. To contrast with this two-stage approach, we refer to traditional gradient-based NAS approaches as single stage methods.
@@
While to date, weight-entangled spaces have only been explored with two-stage methods, and cell-based spaces have only been optimized with single-stage approaches, in this paper we bridge the gap between these parallel sub-communities. We do so by addressing the challenges associated with integrating off-the-shelf single-stage NAS methods with weight-entangled search spaces.
After a discussion of related work (Section 2), we make the following main contributions:
@@
  • We propose a generalized scheme to apply single-stage methods to weight-entangled spaces, while maintaining search efficiency and efficacy at larger scales (Section 3, with visualizations in Figure 1(c) and Figure 2). We refer to this method as TangleNAS.
@@
  • We propose a fair playground for the comparative evaluation of single and two-stage methods (Section 4.1) and study the effect of weight-entanglement in conventional cell-based search spaces (i.e., NASBench201 and the DARTS search space) (Section 4.2).
@@
  • We evaluate our proposed generalized scheme for single-stage methods across a diverse set of weight-entangled macro search spaces and tasks, from image classification (Section 4.3.1, Section 4.3.2) to language modeling (Section 4.3.3).
@@
  • We conduct a comprehensive evaluation of the properties of single and two-stage approaches including any-time-performance, memory consumption, robustness to training fraction and effect of fine-tuning (Section 5), demonstrating that our generalized gradient-based NAS method achieves the best of single and two-stage methods: the enhanced performance, improved supernet fine-tuning properties, superior any-time performance of single-stage methods, and the low memory consumption of two-stage methods .
To facilitate reproducibility, our code is openly accessible here.
@@ 2 Related Work
Weight-sharing was first introduced in ENAS (Pham et al., 2018), which reduced the computational cost of NAS by 1000$\times$ compared to previous methods. However, since this method used reinforcement learning its performance was quite brittle. Bender et al. (2018) simplified the technique, showing that searching for good architectures is possible by training the supernet directly with stochastic gradient descent. This was followed by DARTS (Liu et al., 2019), which set the cornerstone for efficient and effective gradient-based, single-stage, NAS approaches.
DARTS, however, had prominent failure modes, such as its discretization gap and convergence towards parameter-free operations, as outlined in Robust-DARTS (Zela et al., 2020). Numerous gradient-based one-shot optimization techniques were developed since then (Cai et al., 2019; Nayman et al., 2019; Xu et al., 2019; Dong and Yang, 2019; Hu et al., 2020; Li et al., 2021; Chen et al., 2021b; Zhang et al., 2021). Among these, we highlight DrNAS (Chen et al., 2021b), which we will use in our experiments as a representative of gradient-based NAS methods. DrNAS treats one-shot search as a distribution learning problem, where the parameters of a Dirichlet distribution over architectural parameters are learned to identify promising regions of the search space.
@@
Despite the remarkable performance of single-stage methods, they are not directly applicable to some real-world architectural domains, such as transformers, because of the macro-level structure of these search spaces. DASH (Shen et al., 2022) employs a DARTS-like methodology to optimize CNN topologies (i.e. kernel size, dilation) for a diverse set of tasks, reducing computational complexity by appropriately padding and summing kernels with different sizes and dilations. FBNet-v2 (Wan et al., 2020) makes an attempt along these lines for CNN topologies, but its methodology is not easily extendable to search spaces like transformers with multiple interacting modalities, such as embedding dimension, number of heads, expansion ratio, and depth.
@@
Weight-entanglement, on the other hand, provides a more effective way of weight-sharing exclusive to macro-level architectural spaces. The concept of weight-entanglement was developed in slimmable networks (Yu et al., 2018; Yu and Huang, 2019), OFA (Cai et al., 2020) and BigNAS (Yu et al., 2020) in the context of convolutional networks (see also AtomNAS (Mei et al., 2019)) and later spelled out in AutoFormer (Chen et al., 2021a), where it was applied to the transformer architecture.
@@
Single-path-one-shot (SPOS) methods (Guo et al., 2020) have shown a lot of promise in searching weight-entangled spaces. SPOS trains a supernet by uniformly sampling single paths (one at a time to limit memory consumption) and then training the weights along that path. The supernet training is followed by a black-box search that uses the performance of the models sampled from the trained supernet as a proxy.
OFA used a similar idea to optimize different dimensions of CNN architectures, such as its depth, width, kernel size and resolution. Additionally, it enforced training of larger to smaller sub-networks sequentially to prevent interference between sub-networks. Subsequently, AutoFormer adopted the SPOS method to optimize a weight-entangled space of transformer architectures.
@@
In this work, we demonstrate how single-stage methods can be applied to macro-level search spaces with weight-entanglement, where one can benefit from the time efficiency and remarkable effectiveness of modern differentiable NAS optimizers, while concurrently maintaining memory efficiency of the weight-entangled space. While we chose to adopt DrNAS as the primary approach for our exploration of the weight-entangled space in this study, it is worth highlighting that our methodology is generally applicable to other gradient-based NAS methods.
@@ 3 Methodology: Single-Stage NAS with Weight-Superposition
Our primary goal in this work is to effectively apply single-stage NAS to search spaces with macro-level architectural choices. To reduce the memory consumption and preserve computational efficiency, we propose two major modifications to single-stage methods. In our framework, the operators on any given edge inherit the weights of the largest operator on that edge. This reduces the size of the supernet to the size of the largest individual architecture in the search space.
Figure 1: (a) Two-Stage NAS with WE (Algorithm 3): dotted paths show operation choices not sampled at the given step (b) Single-Stage NAS with WS (Algorithm 4): every operation choice is evaluated independently and contributes to the output feature map with corresponding architecture parameters (c) Single-Stage NAS with WE (Algorithm 1): operation choices superimposed with corresponding arhictecture parameters.
Figure 2: Supernet weight matrix (LHS) adapted to gradient-based methods (RHS). All operation choices are zero-padded to match the dimension of the largest operation and superimposed with given architecture parameter. (a) Weight superposition for a single dimension choice (e.g. kernel size) with architecture parameter $\alpha_{i}$ (b) Weight superposition for combination of multiple dimensions (e.g. embedding dimension and expansion ratio) with architecture parameters $\alpha_{i}$, $\beta_{j}$. White areas inside the dashed gray boundaries indicate zero-padding and color values reflect scaling.
@@
2:  $Cell_{j}\leftarrow DAG(\mathcal{O}_{j},{\mathcal{W}_{max}}_{j})$ /* defined for j=1...M */
3:  $Supernet\leftarrow\cup_{j}^{M}Cell_{j}\cup\mathcal{A}$
@@
12:   $\mathcal{\mathcal{missing}}{W}_{\max}=\mathcal{W}_{\max}-\eta\nabla_{\mathcal{W}_{\max}}\mathcal{L}_{train}(\mathcal{W}_{\max},\mathcal{A})$
@@
Then, instead of weighting the outputs of operations on a given edge by architectural parameters, as single-stage NAS methods do (see Figure 1(b)), we simply take each operation choice, zero-pad that to match the dimensions of the largest operation and sum these terms, each weighted by the corresponding architectural parameter. Figure 2(a) provides an overview of the idea for a single architectural choice such as the kernel size. This is equivalent to taking the largest operation and re-scaling the weights of each sub-operation by the sum of architectural parameters of operations it is embedded in (see right-most weight matrix in Figure 2(a).
This process results in a superposition of operation weights whose structural property is similar to the largest operation. Consequently, a single forward pass suffices to capture the effect of all operational choices, thus making our approach computationally efficient. Furthermore, to accommodate for operations depending on two or more architectural dimension choices we introduce the combi-superposition outlined in Figure 2(b) and Algorithm 2. This allows us to apply any arbitrary gradient-based NAS method, such as DARTS, GDAS or DrNAS, to macro-level search spaces that leverage weight-entanglement, without additional memory consumption and computational cost during forward propagation. We refer to our memory and compute optimal single-stage architecture search method as TangleNAS and provide an overview of the approach in Algorithm 1. The operation $f$ in Algorithm 1 determines the differentiable optimizer used in the method. A $softmax$ function corresponds to DARTS, while sampling from the Dirichlet distribution corresponds to DrNAS. In all experiments of the upcoming sections we use DrNAS as a basic gradient-based NAS method. We therefore simply refer to this combination as TangleNAS throughout the paper.
@@ 4.1 Toy search spaces
We begin the evaluation of TangleNAS on two relatively compact toy search spaces that we designed as a contribution to the community to allow faster iterations of algorithm development:
  • Toy cell space: a small version of the DARTS space; architectures are evaluated on the Fashion-MNIST dataset.
  • Toy conv-macro space: a small space inspired by MobileNet, including kernel sizes and the number of channels in each convolution layer; architectures are evaluated on CIFAR-10.
We describe these spaces in Appendix C, including links to code for these open source toy benchmarks. The results of these experiments are summarized in Tables 2 and 2. Across both of these search spaces, TangleNAS outperforms its two-stage counterparts. Also, DrNAS without weight entanglement performs very poorly on the macro level search space.
Table 1: Evaluation on the toy cell-based search space on the Fashion-MNIST dataset
Search Type Optimizer Supernet type Test acc Search Time(hrs)
Single-Stage DrNAS WS $91.190\pm 0.049$ 6.277
TangleNAS WE $\mathbf{91.300\pm 0.023}$ 6.222
Two-Stage SPOS+RS WE $90.680\pm 0.253$ 15.611
SPOS+ES WE $90.317\pm 0.223$ 13.244
@@
Search Type Optimizer Supernet type Test acc Search Time(hrs)
Single-Stage DrNAS WS $10\pm 0.00$ 12.361
TangleNAS WE $\mathbf{82.495\pm.0.461}$ 8.55
Two-Stage SPOS+RS WE $81.253\pm 0.6717$ 21.676
SPOS+ES WE $81.890\pm 0.800$ 26.359
@@
Table 1: Evaluation on the toy cell-based search space on the Fashion-MNIST dataset
Table 2: Evaluation on the toy conv-macro search space on the CIFAR10 dataset. Note that DrNAS with weight sharing converges to a degenerate architecture here.

### v2

@@ Abstract
Weight sharing is a fundamental concept in neural architecture search (NAS), enabling gradient-based methods to explore cell-based architectural spaces significantly faster than traditional black-box approaches. In parallel, weight-entanglement has emerged as a technique for more intricate parameter sharing amongst macro-architectural spaces. Since weight-entanglement is not directly compatible with gradient-based NAS methods, these two paradigms have largely developed independently in parallel sub-communities. This paper aims to bridge the gap between these sub-communities by proposing a novel scheme to adapt gradient-based methods for weight-entangled spaces. This enables us to conduct an in-depth comparative assessment and analysis of the performance of gradient-based NAS in weight-entangled search spaces. Our findings reveal that this integration of weight-entanglement and gradient-based NAS brings forth the various benefits of gradient-based methods, while preserving the memory efficiency of weight-entangled spaces. The code for our work is openly accessible here.
@@ 1 Introduction
The concept of weight-sharing in Neural Architecture Search (NAS) was motivated by the need to improve efficiency over that of conventional black-box NAS algorithms, which demand significant computational resources to evaluate individual architectures. Here, weight-sharing (WS) refers to the paradigm by which we represent the search space with a single large supernet, also known as the one-shot model, that subsumes all the candidate architectures in that space. Every edge of this supernet holds all the possible operations that can be assigned to that edge. Importantly, architectures that share a particular operation also share its corresponding operation weights, allowing for efficient simultaneous partial training of an exponential number of subnetworks with each gradient update.
@@
Gradient-based NAS algorithms (or optimizers), such as DARTS (liu-iclr19a), GDAS (dong-cvpr19a) and DrNAS (chen-iclr21a), assign an architectural parameter to every choice of operation on a given edge of the supernet. The output feature maps of these edges are thus an aggregation of the outputs of the individual operations on that edge, weighted by their architectural parameters. These architectural parameters are learned using gradient updates by differentiating through the validation loss. Supernet weights and architecture parameters are therefore trained simultaneously in a bi-level fashion. Once this training phase is complete, the final architecture can be obtained quickly, e.g., by selecting operations with the highest architectural weights on each edge as depicted in Figure 1(b). However, more sophisticated methods have also been explored (wang-iclr21a) for this selection.
@@
While gradient-based NAS methods have primarily been studied for cell-based search spaces, a different class of search spaces focuses on macro-level structural decisions, such as the number of channels in a layer of the supernet, or the number of layers which are stacked to form the supernet. In these spaces, all architectures are subnetworks of the architecture with the largest architectural choices, identical to the supernet in this case. These search spaces share weights more intricately via weight-entanglement (WE) between similar operations on the same edge. An example of this for convolutional layers is that the 9 weights of a 3 $\times$ 3 convolution are a subset of the 25 weights of a 5 $\times$ 5 convolution. This paradigm reduces the memory footprint of the supernet to the size of the largest architecture in the space, unlike WS search spaces, where it increases linearly with the number of operation choices.
@@
In order to efficiently search over such weight-entanglement spaces, two-stage methods first pre-train the supernet and then perform black-box search on it to obtain the final architecture. OFA (cai-iclr20a), SPOS (guo-eccv20a), AutoFormer (chen-iccv21b) and HAT (wang-acl20) are prominent examples of methods that fall into this category. Note that these methods do not employ additional architectural parameters for supernet training or search. They typically train the supernet by randomly sampling subnetworks and training them as depicted in Figure 1(a). The post-hoc black-box search relies on using the performance of subnetworks sampled from the trained supernet as a proxy for true performance on the unseen test set. To contrast with this two-stage approach, we refer to traditional gradient-based NAS approaches as single-stage methods.
Figure 1: (a) Two-Stage NAS with WE (Algorithm 3): dotted paths show operation choices not sampled at the given step (b) Single-Stage NAS with WS (Algorithm 4): every operation choice is evaluated independently and contributes to the output feature map with corresponding architecture parameters (c) Single-Stage NAS with WE (Algorithm 1): operation choices superimposed with corresponding architecture parameters. The architecture parameters for the three operation choices are represented by ${[\alpha_{i}]}_{i=1}^{3}$ and ${[\beta_{i}]}_{i=1}^{3}$. The operation weights, or choices, are symbolized by cubes (for convolutional layers) or rectangles (for feedforward layers) in various colors. In scenarios (a) and (c), due to weight entanglement, the smaller weights are effectively structured subsets of the larger weights. Conversely, in (b), through weight-sharing, operation weights are maintained independently from each other. In both (b) and (c), to determine the optimal architecture, the operations associated with the highest architecture parameter value are selected. This selection process applies to the choice of kernel size and the output dimension of the feedforward network.
@@
To date, weight-entangled spaces have only been explored with two-stage methods, and cell-based spaces only with single-stage approaches. In this paper, we bridge the gap between these parallel sub-communities. We do so by addressing the challenges associated with integrating off-the-shelf single-stage NAS methods with weight-entangled search spaces. After a discussion of related work (Section 2), we make the following main contributions:
  1. 1.
@@
    We propose a generalized scheme to apply single-stage methods to weight-entangled spaces while maintaining search efficiency and efficacy at larger scales (Section 3, with visualizations in Figure 1(c) and Figure 2 and Figure 3). We refer to this method as TangleNAS.
  2. 2.
@@
    We propose a unified evaluation framework for the comparative evaluation of single and two-stage methods (Section 4.1) and study the effect of weight-entanglement in conventional cell-based search spaces (i.e., NAS-Bench-201 and the DARTS search space) (Section 4.2).
  3. 3.
@@
    We evaluate our proposed generalized scheme for single-stage methods across a diverse set of weight-entangled macro search spaces and tasks, from image classification (Section 4.3.1 and Section 4.3.2) to language modeling (Section 4.3.3).
  4. 4.
@@
    We conduct a comprehensive evaluation of the properties of single and two-stage approaches (Section 5), demonstrating that our generalized gradient-based NAS method achieves the best of single and two-stage methods: the enhanced performance, improved supernet fine-tuning properties, superior any-time performance of single-stage methods, and low memory consumption of two-stage methods. To facilitate reproducibility, our code is openly accessible here.
@@ 2 Related Work
[…]
@@
DARTS, however, had prominent failure modes, such as its discretization gap and convergence towards parameter-free operations (white-arxiv23a), as outlined in Robust-DARTS (zela-iclr20a). Numerous gradient-based one-shot optimization techniques were developed since then (cai-iclr19a; nayman-neurips19a; wang-ijcai21; dong-cvpr19a; hu-cvpr20a; li-iclr21b; chen-iclr21a; zhang-icml21a). Amongst these, we highlight DrNAS (chen-iclr21a), which we will use in our experiments as a representative of gradient-based NAS methods. DrNAS treats one-shot search as a distribution learning problem, where the parameters of a Dirichlet distribution over architectural parameters are learned to identify promising regions of the search space. Despite the remarkable performance of single-stage methods, they are not directly applicable to some real-world architectural domains, such as transformers, due to the macro-level structure of these search spaces. DASH (shen-neurips22) employs a DARTS-like method to optimize CNN topologies (i.e., kernel size, dilation) for a diverse set of tasks, reducing computational complexity by appropriately padding and summing kernels with different sizes and dilations. FBNet-v2 (wan-cvpr20) and MergeNAS (wang-ijcai21) make an attempt along these lines for CNN topologies, but their methodology is not easily extendable to search spaces like transformers with multiple interacting modalities, such as embedding dimension, number of heads, expansion ratio, and depth.
@@
Weight-entanglement, on the other hand, provides a more effective way of weight-sharing, exclusive to macro-level architectural spaces. In weight-entanglement, operations with weights of smaller dimensionality are structured subsets of the largest dimension. Hence, the total number of parameters in the supernet is the same as the number of parameters in the largest architecture in the space. The concept of weight-entanglement was developed in slimmable networks (yu-iclr18; yu-cvpr19), OFA (cai-iclr20a) and BigNAS (yu-eccv20) in the context of convolutional networks (see also AtomNAS (mei-iclr19)) and later spelled out in AutoFormer (chen-iccv21b) and applied to the transformer architecture.
@@
Single-path-one-shot (SPOS) methods (guo-eccv20a) have shown a lot of promise in searching weight-entangled spaces. SPOS trains a supernet by uniformly sampling paths (one at a time to limit memory consumption) and then training the weights along that path. The supernet training is followed by a black-box search that uses the performance of the models sampled from the trained supernet as a proxy. OFA used a similar idea to optimize different dimensions of CNN architectures, such as its depth, width, kernel size, and resolution. Additionally, it enforced training of larger to smaller subnetworks sequentially to prevent interference between subnetworks. Subsequently, AutoFormer adopted the SPOS method to optimize a weight-entangled space of transformer architectures.
@@
In this work, we demonstrate the application of single-stage methods to macro-level search spaces with weight-entanglement. This approach leverages the time efficiency and effectiveness of modern differentiable NAS optimizers, while maintaining the memory efficiency inherent to the weight-entangled space. Although DrNAS is our primary method for exploring weight-entangled spaces, our methodology is broadly applicable to other gradient-based NAS methods.
@@ 3 Methodology: Single-Stage NAS with Weight-Superposition
Figure 2: Weight superposition with architecture parameters ${\alpha_{i}}_{i=1}^{3}$ for kernel size search. Supernet weight matrix (LHS) is adapted to gradient-based methods (RHS).
Computational Efficiency
Our primary goal in this work is to effectively apply single-stage NAS to search spaces with macro-level architectural choices. To reduce the memory consumption and preserve computational efficiency, we propose two major modifications to single-stage methods. Firstly, the weights of the operations on every edge are shared with the corresponding weights of the largest operator on that edge. This reduces the size of the supernet to the size of the largest individual architecture in the search space.
Figure 3: Combi-superposition with parameters $\alpha_{i}\beta_{j}$. Supernet weight matrix (LHS) is adapted to gradient-based methods (RHS).
Secondly, to compute operation mixture, we take each operation choice, zero-pad to match the size of the largest choice, and then sum these, with each one multiplied by its respective architectural parameter. The resulting operation thus represents a mixture of different choices on an edge. This approach, visualized in Figure 1(c), contrasts with single-stage NAS methods, which weigh the outputs of operations on a specific edge using architectural parameters (refer to Figure 1(b)). Figure 2 provides an overview of the idea for a single architectural choice, such as the kernel size. This is equivalent to taking the largest operation and re-scaling the weights of each sub-operation by corresponding architectural parameters followed by summation of the weights (see the right-most weight matrix in Figure 2).
@@
[…]
@@
12: $\mathcal{\mathcal{}}{W}_{\max}=\mathcal{W}_{\max}-\eta\nabla_{\mathcal{W}_{\max}}\mathcal{L}_{train}(\mathcal{W}_{\max},\mathcal{A})$
@@
Weight Superposition
Weight Superposition is defined as a weighted summation of subsets of the largest weight matrix, to obtain weights with structural properties comparable to the largest operation. Consequently, a single forward pass on the superimposed weights suffices to capture the effect of all operational choices, thus making our approach computationally efficient. See Figure 2 for details on weight superposition of the kernel size.
Combi-Superposition
An operation may depend on two or more architectural decisions. Consider, for example, embedding dimension and intermediate MLP dimension (where the latter depends on the former by a searchable multiplicative factor i.e., the MLP ratio). To accommodate this, we introduce combi-superposition outlined in Figure 3 and in Algorithm 2. Combi-superposition superimposes the weights across multiple different architectural dimensions, allowing for search across different interacting architectural modalities.
Algorithm
These concepts allow us to apply any arbitrary gradient-based NAS method, such as DARTS, GDAS, or DrNAS, to macro-level search spaces that leverage weight-entanglement without incurring additional memory and computational costs during forward propagation. We name this single-stage architecture search method TangleNAS. See Algorithm 1 for an overview of the bi-level optimization framework with weight-superposition. The operation $f$ in Algorithm 1 determines the differentiable optimizer used in the method. For example, $f$ is a softmax function for DARTS and a function that samples from the Dirichlet distribution for DrNAS. We use DrNAS as the primary gradient-based NAS method in all our experiments, and refer to DrNAS used in conjunction with weight-entangled spaces as TangleNAS in the remainder of the paper.
@@ 4.1 Toy search spaces
We begin the evaluation of TangleNAS on two compact toy search spaces that we designed as a contribution to the community to allow faster iterations of algorithm development:
  • Toy cell space: a small version of the DARTS space; architectures are evaluated on the Fashion-MNIST dataset (xiao2017fashion).
  • Toy conv-macro space: a small space inspired by MobileNet, including kernel sizes and the number of channels in each convolution layer as architectural decisions. The architectures are evaluated on CIFAR-10.
We describe these spaces in Appendix D, including links to code for these open source toy benchmarks. The results of these experiments are summarized in Tables 2 and 2. In both of these search spaces, TangleNAS outperforms its two-stage counterparts over 4 seeds. Additionally, DrNAS without weight-entanglement performs extremely poorly on the macro level search space (equivalent to a random classifier).
Search Type Optimizer Supernet type Test acc (%) Search Time (hrs)
Single-Stage DrNAS WS $91.190\pm 0.049$ 6.3
TangleNAS WE $\mathbf{91.300\pm 0.023}$ 6.2
Two-Stage SPOS+RS WE $90.680\pm 0.253$ 15.6
WE $90.317\pm 0.223$ 13.2
@@
Table 1: Evaluation on the toy cell-based search space on the Fashion-MNIST dataset.
Search Type Optimizer Supernet type Test acc (%) Search Time (hrs)
Single-Stage DrNAS WS $10\pm 0.000$ 12.4
TangleNAS WE $\mathbf{82.495\pm 0.461}$ 8.6
Two-Stage SPOS+RS WE $81.253\pm 0.672$ 21.7
WE $81.890\pm 0.800$ 26.4
@@
Table 2: Evaluation on the toy conv-macro search space on the CIFAR-10 dataset.